OpenStudy (anonymous):

Find the inverses of f(x) = 1/8 (x-3) and g(x) = x^3 then find: f^-1(g^-1(1)) g^-1(f^-1(-3)) f^-1(f^-1(6)) g^-1(g^-1(-4)) f(g)^-1 g^-1(f^-1) Pre-calculus just gets complicated everyday..

OpenStudy (anonymous):

My inverses: \[f^{-1}(x) = 8(x + 3)\] \[g^{-1} = \sqrt[3]{x}\]

OpenStudy (anonymous):

please just help me with the first one and i'll do the rest.